OpenStudy (anonymous):

You go out to dinner and run up a bill of $52. You had a 10 percent off coupon that you used. You were also charged a 9 percent tax and left a 17 percent tip. What was your total bill for dinner?

OpenStudy (anonymous):

@rajathsbhat

OpenStudy (anonymous):

@KAOS|abdullah

OpenStudy (anonymous):

I just want to check; I got 59.68. Is that right.?

OpenStudy (anonymous):

i'll calculate and check.....

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

i seem to be getting a different answer

OpenStudy (anonymous):

Really? What? Would you like to see what I did?

OpenStudy (anonymous):

oh....frick wait....i took 54 dollars

OpenStudy (anonymous):

$52-10%=46.80+9%=51.01+17% tip = $59.68

OpenStudy (anonymous):

yup that is the basic calculation

OpenStudy (anonymous):

So. \[\sqrt{} \]

OpenStudy (anonymous):

if u calculated ryt that is the correct answer

OpenStudy (anonymous):

Great.

OpenStudy (anonymous):

thanks for checking.

OpenStudy (anonymous):

no prob.....i'll be going now....got any more problems just tag me...i'll be back sometime

OpenStudy (anonymous):

Okay.